OpenStudy (anonymous):

factor completely 36x^2 + 60x + 25

OpenStudy (luigi0210):

If you look at the front term and last term, you can see that they have different factors. 36 is 6*6, 25=5*5. If you multiply 5*6=30, and 30*2=60, so: \[\LARGE (6x+5)(6x+5)~~or~~(6x+5)^2\]